OpenStudy (anonymous):

log3 5x + 6 = 4 I am trying to solve for x

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

Solving the equation would be log(3)= exponent 5 sqrt -2. But i dont know if thats solving for x

OpenStudy (whpalmer4):

Is the log of 5x, or 5x+6?

OpenStudy (whpalmer4):

\[\log_3 (5x+6) = 4\]or \[\log_3(5x)+6=4\]

OpenStudy (anonymous):

its log3 (5x+6) = 4

OpenStudy (whpalmer4):

Okay. \[\log_3 (5x+6) = 4\]That means that \[3^4 = 5x+6\]because \[\log_b u = a\]\[b^a=u\]

OpenStudy (whpalmer4):

also \[b^{\log_b u} = u\]so\[3^{\log_3(5x+6)} = 3^4 = 5x+6\]

OpenStudy (anonymous):

okay thank you. with this equation i am receiving an answer of 15

OpenStudy (whpalmer4):

that's correct. if you do the other problem, the answer is \(\dfrac{1}{4 5}\)

OpenStudy (anonymous):

thank you